OpenStudy (anonymous):

A rectangle has an area of 40 square units. The length is 6 units greater than the width. What are the dimensions of the rectangle? 8 by 5 10 by 4 11 by 9 13 by 7

OpenStudy (anonymous):

@ganeshie8 @texaschic101 @triciaal @freckles @Nnesha @Michele_Laino

OpenStudy (anonymous):

40 = L*W L = W+6 Substitute. 40 = (W+6)*W

OpenStudy (anonymous):

40 = W^2 + 6W 0 = W^2+6W-40

OpenStudy (anonymous):

Then you factor, so 0 = (W-4)(W+10) So, the width is either |4| or |-10| That means that the length is either 10 or 4, by plugging it into 40=L*W and solving for L So 10 x 4

OpenStudy (anonymous):

Thanks